### v1

# Towards Agentic AI on Particle Accelerators

Abstract
As particle accelerators grow in complexity, traditional control methods face increasing challenges in achieving optimal performance. This paper envisions a paradigm shift: a decentralized multi-agent framework for accelerator control, powered by Large Language Models (LLMs) and distributed among autonomous agents. We present a proposition of a self-improving decentralized system where intelligent agents handle high-level tasks and communication and each agent is specialized control individual accelerator components.
This approach raises some questions: What are the future applications of AI in particle accelerators? How can we implement an autonomous complex system such as a particle accelerator where agents gradually improve through experience and human feedback? What are the implications of integrating a human-in-the-loop component for labeling operational data and providing expert guidance? We show two examples, where we demonstrate viability of such architecture.
1 Introduction
Particle accelerator operation involves a complex interplay of interrelated systems, each requiring precise control and optimization. Traditionally, these systems have been managed through a combination of human expertise and highly specialized algorithms with recent developments including machine learning techniques [7] where some are successfully deployed in operation like e.g[11, 8, 29].
Most of these algorithms are designed to operate in a narrow domain of accelerator functionality. The systems have noise, tight operational limits, and deterministic interplay through the control system to ensure human and equipment safety, preventing scenarios that could lead to hazardous conditions or equipment damage. While effective, this approach can sometimes struggle to achieve optimal overall performance due to the challenges of integrating these disparate systems. Even when a sufficiently comprehensive model of the facility is constructed, it can be highly sensitive and prone to problems as the accelerator transitions through different states or experiences drifts and thus requires human intervention for setup and ongoing maintenance.
This paper proposes a paradigm shift in accelerator control systems, envisioning a future where self-improving agents control separated sub-components using algorithms that inform operators of significant events. In this framework, we show how autonomous agents controlled by LLMs serve not only in the creation and retrieval of accelerator operation documentation but also as high-level coordinators, handling complex tasks such as the interpretation of the machine state from analyzing control system channels or inter-agent communication with a degree of autonomy in decision-making and communication. Meanwhile, dedicated agents controlled by LLMs execute specific, well-established calculations or operations within their domains of expertise. This decentralized approach, illustrated in our conceptual diagram Fig. 1, showcases a potential future for the operation of particle accelerators with the flexibility to deploy more advanced agents.
By exploring this forward-looking perspective, we aim to stimulate discussion on the future of particle accelerator operations and the broader implications of decentralized AI architectures in scientific instrumentation. The proposed framework offers potential solutions to the challenges of integrating disparate systems and managing complex state transitions, while also providing a flexible architecture that can adapt to evolving requirements and technological advancements.
Figure 1: Schematic overview of a decentralized agent-based control architecture. The left diagram illustrates the modular structure where each agent, consisting of devices, algorithms, and agent controllers, manages specific subsystems. The right figure shows these agents interact with both the physical components of the accelerator and the control room, enabling prompting current state, monitoring, decision-making, and adaptive responses to complex operational scenarios.
2 Related Work
Machine learning techniques have increasingly been applied to various aspects of accelerator physics in recent years. Edelen et al[7] provide a comprehensive review of machine learning applications in particle accelerators, covering areas such as beam diagnostics, control systems, and performance optimization.
2.1 Agentic AI and LLMs in Complex Reasoning Tasks
There is a rise of agentic AI that has demonstrated remarkable capabilities in complex reasoning tasks [2]. According to [27], an effective agent in an agentic system should possess autonomy to operate independently, reactivity to respond to environmental changes, and proactiveness to take initiative in pursuing its objectives, enabling it to function intelligently and adaptively in complex environments. We will highlight the most relevant works that align with our problem.
Yao et al[28] introduced the ReAct framework, which showed how LLMs can effectively combine reasoning and acting in multi-step tasks and is widely used to operate in complex multi-step reasoning.
The field of use of AI agents to effectively engage with their environment and complete a wide array of tasks has been gaining more and more popularity. Park et al[19] introduce generative agents that use LLMs to simulate human behavior in interactive environments, showcasing how LLMs can be utilized to model intricate multi-agent interactions and environmental complexities. Wang et al[25] uses an LLM-powered agent to autonomously explore, acquire skills, and make discoveries in unfamiliar environments using components like an automatic curriculum and skill library. AgentVerse [3] shows a framework for collaborative groups of LLM-powered expert agents, showing improved performance over single agents in various tasks. This multi-agent approach leverages LLMs for high-level tasks while enabling specialized components to work together autonomously on complex problems. AutoGen [26] framework allows building LLM applications using multiple conversable agents. They show that autonomous, communicating agents can control subcomponents of a system. Particle accelerator control often utilizes operation sequencing [9, 1], there LLM+P [17] can help by translating natural language into formal planning languages, potentially bridging the gap between planning, natural language, and accelerator controls. To deal with environment changes, Shinn et al[23] proposes Reflexion, a framework that reinforces language agents through linguistic feedback rather than traditional weight updates. Reflexion agents verbally reflect on task feedback and maintain reflective text in an episodic memory buffer to improve decision-making in subsequent trials. To impose the factuality, [6] proposes multiagent debate where multiple LLMs debate to improve reasoning and factual accuracy of results which is essential for reliable operations. LLMs have gained prominence as coding assistants [14] demonstrating the capability to function as coders [21], problem solvers [22] or even data scientists [16]. Automated Design of Agentic Systems [13] aims to automatically create AI agent designs by having a meta-agent iteratively program new agents in code.
2.2 LLMs as Assistants in Operations
Carrasco et al[2] explores fine-tuned LLMs for autonomous spacecraft control in simulations, showing their efficacy in handling language-based inputs and outputs. It shows the potential of using LLMs for complex tasks like accelerator control.
In accelerators, Mayet’s [18] GAIA system uses LLMs with the ReAct framework [28] to assist in operations by integrating multiple expert tools e.g. knowledge retrieval, machine control, and Python script generation for autonomous and semi-autonomous management of complex accelerator environments.
In [24] they show joint efforts across particle accelerator facilities to utilize Retrieval Augmented Generation (RAG) models and other AI techniques for enhancing eLogs usability and automation, aiming to unlock operational insights and improve data accessibility.
3 Features of the System
Four key aspects of agents can significantly enhance operational effectiveness. First, operating particle accelerators is complex and requires extensive human expertise. We suggest that agents can gradually improve through experience [23] where the system incorporates continuous learning from operational data and learns based on outcomes or from "human-in-the-loop". It can be an important step forward since most of these particle accelerators operate in already high reliabilities above 90%.
Second, the potential to uncover causal relationships in accelerator operations by adding Reasoning Agents [10]. LLM-powered agents can provide a human-readable interface to complex machine operations, enabling faster exploration of causal connections via reasoning about the information available (or gradually reasoning about it via ReAct [28]). By integrating agents designed to reveal casual relationships, we aim to enhance system interpretability and streamline diagnostics. This approach could reduce the learning curve for new operators and offer valuable insights to experienced physicists, leading to more efficient accelerator management.
The third key component is agent autonomy implementable via rules, decision trees, or language models. LLMs are currently preferred due to their natural language interface and intelligent decision-making. LLM high computational needs may cause delays, limiting use in real-time control. The choice of autonomy implementation thus requires balancing LLMs’ flexibility against performance needs in time-sensitive applications.
Lastly, other interesting examples of agents can be: planning agent [17] to execute and run complex plans with standard tools, such as to execute tasks at e.g. European XFEL defined in Taskomat [9] or [1] at Fermi, coding agents [21], and data scientist agents [16] to summarize.
4 Examples
In this section, we present two examples, where the aforementioned features can be directly applicable by integrating them into the control systems like [12, 4, 20] and enhancing operational efficiency and stability: the Advanced Light Source (ALS) orbit feedback system and the European XFEL longitudinal feedback manager.
4.1 ALS Example: Orbit Feedback
Maintaining precise control over the beam orbit through orbit feedback systems is crucial for the stable operation of a storage ring. However, this task can be complicated in practice by machine drifts caused by environmental factors or by maintenance activities that alter the operational characteristics of specific sections of the ring. Under the current operational paradigm, when the orbit feedback system fails to converge during user operation, a physicist must intervene. This process typically involves a detailed analysis of the situation, comparing current conditions with previous runs, and relying heavily on experience to make a judgment call on how to adjust the system for optimal performance.
As illustrated in Fig. 2, in our envisioned framework, this diagnostic role would be managed by a specialized feedback agent. Upon detecting abnormal behavior, the agent would identify the underperforming area and consult the logbook agent for recent events that could explain the issue. For instance, if maintenance work had occurred in the affected sector, the logbook agent would provide this information. The feedback agent would then draft a report outlining the problem, its likely root cause, and a suggested course of action, which would be sent to the control room for review.
Figure 2: Diagram of the orbit feedback agent. The Orbit FB Agent detects abnormal behavior, consults the Logbook Agent for relevant events, and suggests action for the control room to review.
4.2 European XFEL Longitudinal Feedback Manager
The particle beam in a linear accelerator, like the European XFEL, is stabilized within well-defined, narrow parameter spaces through the use of various feedback loops. To effectively orchestrate the complex interplay of these feedback loops, expert systems are often employed [15, 5], which encapsulate the logic of this interaction in software. A Feedback Agent can be trained to learn the possible and desired states, illustrated in Fig. 3. In the next step, such an agent would initially be able to assist the operator as a kind of recommender system and, potentially, could take over the entire operation of all longitudinal feedback loops in a later step.
Figure 3: Diagram of the Longitudinal feedback manager. Left: The operator sets feedback loops and the FB Manager Agent and Logbook Agent observe and recommend actions to the operator. Right: FB Manager Agent takes over the operation and just informs the operator.
Examples of additional agents can be: Orbit Feedback Agent for controlling transversal beam parameters; Undulator Agents for optimizing photon beam production; Beamline Agents overseeing experimental stations; and Feedback Coordination Agent acting as a central coordinator to orchestrate overall feedback operation. Each agent functions as a recommender system, assisting human operators in decision-making. As confidence in their performance grows, they could potentially take over the entire operation of individual feedback, e.g. through [23]. Crucially, these agents maintain constant communication with each other. Before implementing any changes, each agent informs the others of its intended actions. This allows for a collective evaluation of the proposed changes, ensuring that they don’t lead to suboptimal or dangerous configurations.
5 Conclusion
This paper proposes a paradigm shift in particle accelerator control through a decentralized multi-agent framework powered by LLMs. By integrating AI agents for high-level tasks and specialized agents for component management, we address the increasing complexity of modern accelerator systems.
Two examples demonstrate the potential of AI agents in assisting complex accelerator tasks, opening the way for higher autonomy. While this approach offers exciting possibilities for improved performance, our practical examples support a vision for more intelligent and adaptive accelerator operations.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately reflect the paper’s contributions and scope. They present the main idea of a decentralized multi-agent framework for accelerator control powered by Large Language Models (LLMs). The introduction outlines the challenges in current accelerator control systems and proposes the paradigm shift, which aligns with the paper’s overall content and examples provided.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: It primarily focuses on presenting a conceptual framework and discussing potential applications of LLM-powered agents in particle accelerator control, but briefly mentions potentially main issue with LLM latency.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper does not include formal theoretical results or proofs. It primarily focuses on presenting a conceptual framework and discussing potential applications of LLM-powered agents in particle accelerator control.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [No]
  16. Justification: The paper does not present experimental results. Instead, it proposes a conceptual framework and provides examples of how it could be applied in specific accelerator control scenarios. As such, there are no experimental results to reproduce.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [No]
  20. Justification: The paper does not mention any experiments requiring code or data. It is a conceptual paper proposing a framework for future implementation rather than presenting implemented systems or experiments.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [No]
  24. Justification: The paper does not include experiments. It presents a conceptual framework and potential applications without conducting specific experiments or providing experimental details.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: The paper does not include experiments or statistical analyses. It is a conceptual paper proposing a framework for future implementation.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [No]
  32. Justification: The paper does not include experiments or discuss compute resources needed for implementations. It is a conceptual paper proposing a framework for future work.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: The paper discusses potential positive and negative societal impacts, which is in line with ethical considerations in AI research.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: It mentions the potential for improved accelerator performance and efficiency, while also acknowledging potential risks such as the need for human oversight and the importance of safety considerations in autonomous systems.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [No]
  44. Justification: The paper does not release any data or models that have a high risk for misuse. It is a conceptual paper proposing a framework for future implementation rather than releasing actual systems or data.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: The paper cites and credits various existing works and frameworks throughout the text, particularly in the Related Work section. While specific license information is not provided for each reference, the proper attribution of ideas and previous work is evident.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [No]
  52. Justification: The paper does not introduce or release new assets such as datasets, code, or models. It proposes a conceptual framework for future implementation.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [No]
  56. Justification: The paper does not involve crowdsourcing or research with human subjects. It focuses on proposing a conceptual framework for AI-powered accelerator control systems.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [No]
  60. Justification: The paper does not involve research with human subjects, so IRB approval or equivalent is not applicable to this work.